# A Survey of Cross-domain Graph Learning: Progress and Future Directions

### v1

@@ Abstract
Graph learning plays a vital role in mining and analyzing complex relationships involved in graph data, which is widely used in many real-world applications like transaction networks and communication networks. Foundation models in CV and NLP have shown powerful cross-domain capabilities that are also significant in graph domains. However, existing graph learning approaches struggle with cross-domain tasks. Inspired by successes in CV and NLP, cross-domain graph learning has once again become a focal point of attention to realizing true graph foundation models. In this survey, we present a comprehensive review and analysis of existing works on cross-domain graph learning. Concretely, we first propose a new taxonomy, categorizing existing approaches based on the learned cross-domain information: structure, feature, and structure-feature mixture. Next, we systematically survey representative methods in these categories. Finally, we discuss the remaining limitations of existing studies and highlight promising avenues for future research. Relevant papers are summarized and will be consistently updated at: https://github.com/cshhzhao/Awesome-Cross-Domain-Graph-Learning.
Index Terms:
Graph Learning, Cross-domain, Structure, Feature, Structure-feature, Taxonomy.
@@ I Introduction
Graph, or graph theory, has been playing an increasingly crucial role across numerous fields, actively contributing to various industries in the modern world, including recommendation [1, 2], finance [3, 4], weather [5], pharmaceuticals [6, 7], and communications [8, 9]. Graph data can model the structural characteristics between nodes, offering significant advantages in capturing interactions and dependencies across structures and features. Many real-world datasets—such as citation networks [10], molecular [7], and transportation networks [11]—are naturally represented as graphs. To efficiently process and analyze these complex graph data, Graph Neural Networks (GNNs) [12, 13, 14] have emerged as a leading technique. The primary goal of GNNs is to learn the relationships between nodes and their neighbors through recursive message passing and aggregation mechanisms, thereby obtaining expressive representations at the node, edge, or graph level for various downstream tasks.
@@
Figure 1: The purpose of cross-domain graph learning aims to integrate graph knowledge across a myriad of graph domains and transfer the knowledge to other domains.
@@
In recent years, the rapid development of foundation models, especially the large language models (LLMs), has significantly revolutionized the way people live and work, such as ChatGPT in natural language processing (NLP) and DALL-E 2 in computer vision (CV). Compared to previous deep learning models, the core advantage of foundation models lies in their ability to train on data from diverse domains, subsequently transferring the learned knowledge across different domains, encompassing three key capabilities: cross-modal, cross-domain, and cross-task [15, 16]. These three capabilities are also crucial in graph domains. For instance, cross-modal capabilities allow social networks to combine user relationships (graph structure) with textual content (text modality) to more accurately predict user interests, thus enhancing recommendation systems. Cross-domain capabilities facilitate knowledge transfer between different domains—such as leveraging insights from transportation networks to optimize power load management in power grid networks, or using rich user interaction data from social networks to improve anomaly detection in banking transaction networks with limited data. Cross-task capabilities enable models trained for one task (e.g., node classification to predict a paper’s research field) to be directly applied to another task (e.g., link prediction to identify paper citations), thus enhancing adaptability and efficiency across tasks.
@@
The importance of cross-domain graph learning (CDGL). In graph domains, the key to achieving this vision lies in finding an effective way to integrate cross-domain knowledge that allows for straightforward generalization across domains. In detail, single-domain cross-modal problems (e.g., integrating images, text, and graphs) can benefit from advancements in CV and NLP, where many methods have successfully leveraged multi-modal information [17]. Additionally, single-domain cross-task problems can be addressed using classical graph pretraining models [18, 19, 20]. However, when dealing with cross-domain problems, the heterogeneity and diversity of graph data make it challenging for data from a single domain to capture a universal graph representation. Figure 1 shows the overview of CDGL, which can help better understand the differences across multi-domain graph data [21]:
@@
    Structural differences: Graphs from different domains vary in node connectivity, with some being densely connected and others sparse. For instance, the social network among hotel staff forms a structured and top-down hierarchy with sparse connections, and the citation network among academic papers is densely interconnected.
@@
    Feature differences: The dimensions and semantics of node features also differ across domains, such as social networks and molecular networks.
This diversity poses challenges for effectively integrating cross-domain graph information and producing universal cross-domain graph representations. To this end, CDGL should be considered the crucial step towards graph foundation models. Through CDGL, commonalities can be extracted from graphs across different domains, enabling the learning of graph representations with broad transferability.
[…]
@@
[…]
@@
Motivations. Although CDGL has increasingly received attention in graph domains—especially due to its role in constructing graph foundation models—this rapidly expanding field still lacks a systematic review, which hinders researchers from gaining a comprehensive understanding of the field and exploring future directions. Concretely, previous graph-based cross-domain methods were usually designed for specific scenarios, with limited cross-domain scope and lacking universality. Therefore, they are mainly included as subcategories in surveys within particular domains, such as graph neural network-based methods included in recommendation system surveys [1, 84].
To fill this gap, in this work, we aim to provide a systematic survey on CDGL, helping researchers better understand the current state of research and its challenges. Specifically, we go through existing representative CDGL works to organize our first-level taxonomy, as shown in Figure 2, categorizing based on the information captured by graph neural networks used for cross-domain transfer (i.e., structure-oriented, feature-oriented, and mixture-orineted) throughout the entire model pipeline. We further refine our taxonomy and introduce more granularity to the initial categories.
Contributions. The contributions of this work can be summarized from the following three aspects. (1) A structured taxonomy. A broad overview of the field is presented with a structured taxonomy that categorizes existing works into three categories (Figure 2). (2) A comprehensive review. Based on the proposed taxonomy, the current research progress of cross-domain techniques for graph learning is systematically delineated. To the best of our knowledge, this is the first comprehensive survey for CDGL. (3) Some future directions. We discuss the remaining limitations of existing works and point out possible future directions.
@@ II-A Notations
Consider a graph instance represented as $G=\{V,E\}$, where $V=\{v_{1},v_{2},\ldots,v_{N}\}$ is the set of nodes with $N$ elements. The set $E\subseteq V\times V$ defines the connections between nodes. Associated with each node $v_{i}$ is a feature vector $x_{i}\in\mathbb{R}^{D}$. To describe the connectivity within the graph, we use the adjacency matrix $A\in\{0,1\}^{N\times N}$, where the matrix entry $A_{i,j}$ is defined as 1 if and only if there is an edge $(v_{i},v_{j})\in E$. A graph data space, including a node space and an edge space in a graph space, can be denoted as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ are the set of nodes and edges in the corresponding graph space.
@@ II-B Graph Representation Learning
Shallow Embedding Methods. Shallow embedding methods, such as node2vec [85] and DeepWalk [86], aim to map nodes into low-dimensional embeddings by preserving the network’s similarity structure. These methods utilize techniques like factorization-based methods or random walks to achieve this goal. While providing flexibility for various downstream tasks, shallow embedding methods face limitations: they cannot generate embeddings for unseen nodes and lack the capability to incorporate node features effectively.
Deep Graph Neural Networks (GNNs). Deep GNNs, including Graph Convolutional Networks (GCN) [12] and Graph Attention Networks (GAT) [13], maintain input node features as constants and optimize model parameters for specific tasks. These networks typically employ a message-passing mechanism, where nodes iteratively exchange and aggregate information with their neighbors. This process leads to more expressive graph representations that can effectively handle more complex graph-based tasks. In detail, the forward process of the message-passing mechanism can be defined as:
$h_{i}^{(l)}=\mathbf{U}\left(h_{i}^{(l-1)},\mathbf{M}(\{h_{i}^{(l-1)},h_{j}^{(l-1)}\mid v_{j}\in\mathcal{N}_{i}\})\right)$
where $h_{i}^{(l)}$ is the feature vector of node $v_{i}$ in the $l$-th layer, and $\mathcal{N}_{i}$ is a set of neighbor nodes of node $v_{i}$. $\mathbf{M}$ denotes the message passing function of aggregating neighbor information, $\mathbf{U}$ denotes the update function with central node feature and neighbor node features as input. By stacking multiple layers, GNNs can aggregate messages from higher-order neighbors.
@@ II-C Cross-domain Learning
Definitions. Given the source domains $D_{S}=\{D_{S_{1}},\dots,D_{S_{m}}|D_{S_{i}}=\{X_{S_{i}},Y_{S_{i}}\},i=1,\dots,m\}$ and target domains $D_{T}=\{D_{T_{1}},\dots,D_{T_{n}}|D_{T_{j}}=\{X_{T_{j}},Y_{T_{j}}\},j=1,\dots,n\}$, where for any $i$ and $j$, $|D_{T_{j}}\cap D_{S_{i}}|\geq 0$, the objective of general cross-domain learning is to map the data from source domains to a shared embedding space and then learn a set of task functions based on this shared space for the target domains. Each dataset is referred to as a distinct domain, where a domain is characterized by a unique data distribution and task context.
First, we define the model $M$ as a source encoder that maps all $X_{S_{i}}\in D_{S}$ to a unified embedding space $E$, which can be also applied for target domains:
$M:X_{S_{i}}\rightarrow E,\ \text{for}\ i=1,\dots,m.$ (1)
Next, we define the set of task functions for the target domains $F=\{f_{T_{1}},\dots,f_{T_{n}}\}$, where each task function $f_{T_{j}}$ aims to map the embeddings $M(X_{T_{j}})$ to the corresponding task result $Y_{T_{j}}$:
$F=\{f_{T_{j}}:M(X_{T_{j}})\rightarrow Y_{T_{j}},\ \text{for}\ j=1,\dots,n\}.$ (2)
In this framework, the objective is to learn the model $M$ and the task functions $f_{T_{j}}$ such that the features across all domains are unified into a shared embedding space, allowing effective knowledge transfer across domains, which in turn improves generalization and performance on the target tasks.
The objective function for training $M$ can be defined as:
$\min_{M}\sum_{i=1}^{m}\mathbb{E}_{x\sim P_{S_{i}}(x)}\left[L_{M}(M(x),Y_{S_{i}})\right],$ (3)
where $L_{M}$ can be an unsupervised or supervised loss function (e.g., contrastive loss or classification loss), and $P_{S_{i}}(x)$ represents the distribution of the source domain data.
@@
$\min_{f_{T_{j}}}\sum_{j=1}^{n}\mathbb{E}_{x\sim P_{T_{j}}(x)}\left[L_{T}(f_{T_{j}}(M(x)),Y_{T_{j}})\right],$ (4)
@@
Note that these two processes can be conducted either through joint training or separate training. The conditions $|D_{T_{j}}\cap D_{S_{i}}|$ reflect, to some extent, the difficulty of transferring cross-domain knowledge during the training process.
@@
    NLP. The development of cross-domain learning in NLP has progressed significantly, driven by innovations in model architectures, training strategies, and data availability. Early efforts relied on domain adaptation techniques, such as feature alignment and adversarial learning, to transfer knowledge from resource-rich domains to low-resource ones. Alongside these, domain generalization techniques, which aim to improve model robustness by learning domain-invariant representations, have also been widely explored. The rise of pre-trained language models (PLMs), like BERT and GPT, enabled generalization across domains by leveraging contextual embeddings and task-specific fine-tuning. More recently, large language models (LLMs), such as GPT-3 and GPT-4, have demonstrated strong cross-domain capabilities through massive pretraining on diverse corpora, excelling in zero-shot and few-shot learning. Furthermore, multi-modal techniques have extended cross-domain learning beyond text, allowing models to integrate information from other modalities, such as images and structured data, to address complex tasks like visual question answering and multi-modal dialogue.
@@
    CV. The development of cross-domain learning in computer vision (CV) follows a similar trajectory to that of NLP. Early approaches rely on domain adaptation techniques such as adversarial training, feature alignment, and style transfer to address challenges posed by domains with different visual characteristics. As the field advances, domain generalization methods, including invariant feature learning and meta-learning, emerge to improve model robustness to unseen domains. The introduction of pre-trained vision models like ResNet and Vision Transformers (ViTs) further the progress by enabling models to generalize across domains through the reuse of learned features, which can then be fine-tuned for downstream tasks. Recently, multi-modal models like CLIP and Sora demonstrate exceptional cross-domain capabilities by integrating visual and language understanding. For instance, Sora excels in tasks like image captioning, visual question answering, and zero-shot reasoning, benefiting from large-scale pretraining on multi-modal datasets. Just as in NLP, the combination of visual and textual data allows these models to achieve superior cross-domain performance.

### v2

@@ Abstract
Graph learning plays a vital role in mining and analyzing complex relationships within graph data and has been widely applied to real-world scenarios such as social, citation, and e-commerce networks. Foundation models in computer vision (CV) and natural language processing (NLP) have demonstrated remarkable cross-domain capabilities that are equally significant for graph data. However, existing graph learning approaches often struggle to generalize across domains. Motivated by recent advances in CV and NLP, cross-domain graph learning (CDGL) has gained renewed attention as a promising step toward realizing true graph foundation models. In this survey, we provide a comprehensive review and analysis of existing works on CDGL. We propose a new taxonomy that categorizes existing approaches according to the type of transferable knowledge learned across domains: structure-oriented, feature-oriented, and mixture-oriented. Based on this taxonomy, we systematically summarize representative methods in each category, discuss the key challenges and limitations of current studies, and outline promising directions for future research. A continuously updated collection of related works is available at: https://github.com/cshhzhao/Awesome-Cross-Domain-Graph-Learning.
@@ I Introduction
Graph, or graph theory, has been playing an increasingly crucial role across numerous fields, actively contributing to various industries in the modern world, including recommendation [zang2022relatedsurvey2, liu2020modelling], finance [li2023digaBank1, deprez2024networkBank2], weather [liu2025cirt], pharmaceuticals [li2022Pharm1, li2023Pharm2molecularalign], and communications [ji2022communication2, zhao2023communication1]. Graph data can model the structural characteristics between nodes, offering significant advantages in capturing interactions and dependencies across structures and features. Many real-world datasets (e.g., social, citation, molecular, e-commerce, and transportation networks [sen2008citation1, li2023Pharm2molecularalign, zhu2024graphclip, yu2024temporalEcommerce1, jin2024urbanTransport1]) are naturally represented as graphs. To efficiently process and analyze these complex graph data, Graph Neural Networks (GNNs) [kipf2016gcn, velivckovic2018gat, liusegno] have emerged as a leading technique. The primary goal of GNNs is to learn the relationships between nodes and their neighbors through recursive message passing and aggregation mechanisms, thereby obtaining expressive representations at the node, edge, or graph level for various downstream tasks.
@@
Figure 1: Cross-domain graph learning aims to integrate knowledge from multi-domain source graphs and transfer it to diverse target domains.
@@
In recent years, the rapid development of foundation models, especially the large language models (LLMs) [chang2024llmsurvey1, min2023llmsurvey2], has significantly revolutionized the way people live and work, such as ChatGPT in natural language processing (NLP) [achiam2023gpt4] and DALL-E 3 in computer vision (CV) [betker2023DALLE3]. Compared to previous deep learning models, the core advantage of foundation models lies in their ability to train on data from diverse domains, subsequently transferring the learned knowledge across different domains, encompassing three key capabilities: cross-modal, cross-domain, and cross-task [zhao2023LlmThreeCapabilities1, sun2023LlmThreeCapabilities2]. These three capabilities are also crucial in graph domains. For instance, cross-modal capabilities allow user-item interaction networks to combine user relationships (graph structure) with textual content (text modality) to more accurately predict user interests, thus enhancing recommendation systems [wu2025graphCrossModal]. Cross-domain capabilities facilitate knowledge transfer between different domains—such as leveraging insights from transportation networks to optimize power load management in power grid networks [huang2024towardCrossDomainPower1, yang2025recentCrossDomainPower2], or using rich user interaction data from social networks to improve anomaly detection in banking transaction networks with limited data [li2023digaBank1]. Cross-task capabilities enable models trained for one task (e.g., node classification to predict a paper’s research field) to be directly applied to another task (e.g., link prediction to identify paper citations), thus enhancing adaptability and efficiency across tasks [sun2023crossTask1, wang2025GIT]. Therefore, realizing a graph foundation model (GFM) equipped with the three key capabilities is essential for advancing the field of graph learning.
@@
The importance of cross-domain graph learning (CDGL). To propel graph learning into the era of true foundation models, the top priority is to develop an effective method for distilling commonalities from cross-domain graph data, enabling the representation space of pretrained graph models to generalize effortlessly across diverse domains. In detail, single-domain cross-modal graph learning (e.g., integrating images, text, and graphs) can benefit from advancements in CV and NLP, where many methods have successfully integrated multi-modal information [jin2024crossModalsurvey1]. Additionally, single-domain cross-task graph learning can be addressed using classical graph pretraining models [you2020crossTask2, sun2023crossTask1, zhao2024crossTasksurvey1]. However, when handling cross-domain graph learning (Figure 1), the diversity and complexity of graph data across domains make it difficult for graph models to extract universal knowledge from source domains and transfer it to target domains [zhao2024GCOPE]. This challenge primarily manifests in two aspects:
@@
    Structural differences: Graphs from different domains can vary significantly in both connectivity and scale. Some graphs are sparsely connected, such as citation networks like Cora [he2025cora], where most papers cite only a few others. In contrast, graphs like those in COLLAB [yanardag2015COLLAB] are densely connected, as researchers often collaborate extensively within the same subfield. In terms of scale, taking Flickr [tang2009Flickr, gao2024flickr] as an instance, it forms a large-scale social network graph where users are nodes and “contact” links are edges, whereas molecular datasets consist of many small graphs [li2022molecularsmall], each representing a single molecule. These differences constitute diverse structural patterns; each domain’s unique characteristic can trigger structure-level negative transfer [zhu2021structurelevelNegativeTransfer].
@@
    Feature differences: Node features in graphs from different domains are domain-specific, resulting in differences in both dimensionality and semantics. For example, in citation networks, node features typically represent the title and abstract of a paper, resulting in high-dimensional node features [li2024zerog]. In contrast, in molecular graphs, node features describe properties of atoms (e.g., atom type, atomic mass, hybridization), which are fewer in number, leading to low-dimensional node features [li2022molecularsmall]. Semantic gaps hinder identifying cross-domain feature commonalities [wu2024GDomainAdaptationSurvey2], while dimensional mismatches can even prevent graph models from learning across domains [zhao2024GCOPE].
Taken together, these differences pose significant challenges for effectively learning from cross-domain graph knowledge and producing generalizable graph representations. Drawing on the foregoing analysis, CDGL emerges as the pivotal research direction toward genuine GFMs. By harnessing CDGL, we can distill shared structural and feature patterns across different graph domains, endowing learned representations with robust, domain-agnostic transferability.
@@
Structure-oriented Generative GraphControl [zhu2024graphcontrol], GA2E [hu2024GA2E], OpenGraph [xia2024opengraph], UniAug [tang2024UniAug], GFT [wang2024gft], GFSE [chen2025gfse] Contrastive GCC [qiu2020gcc], PCRec[wang2021pcrec], GRADE [wu2023GRADE], APT [xu2023APT] , FedStar [tan2023FedStar], BooG [cheng2024BooG], ProCom [wu2024procom], RiemannGFM [sun2025RiemannGFM] Feature-oriented In-space KTN [yoon2022KTN], NaP [wang2024NAP] , GPF-plus [fang2024gpfplus], RELIEF [zhu2024relief], DAGPrompT [chen2025DAGPrompT], GADT3 [pirhayatifard2025crossdomainGADT3] Cross-space OFA [liu2024OFA] , GraphAlign [hou2024graphalign], CDFS-GAD [chen2024CDFSGAD], ZeroG [li2024zerog], UNPrompt [niu2024UNPrompt] Mixture-oriented Sequential PTGB [yang2023ptgb], SDGA [qiao2023SDGA], SOGA [mao2024SOGA], GSPT [song2024GSPT], ARC [liu2024ARC], UniGraph [he2024unigraph], AnyGraph [xia2024anygraph], GraphLoRA [yang2024graphlora], SAMGPT [yu2025SAMGPT], MDGFM [wang2025MDGFM], MLDGG [tian2025mldgg] Unified Graph-based UDA-GCN [wu2020UDAGCN], DA-GCN[Guo2021DAGCN], COMMDER[ding2021COMMANDER], GCFL[xie2021gcfl], CDGEncoder [hassani2022CDGEncoder], CrossHG-Meta [zhang2022CrossHGMeta], METABrainC [yang2022METABrainC], CCDR [xie2022CCDR], GAST [zhang2022GAST], DH-GAT[xu2023DHGAT], ALEX[yuan2023alex], ACT [wang2023ACT], CDTC [zhang2023CDTC], DGASN [shen2023DGASN], STGP [hu2024stgp], GCOPE [zhao2024GCOPE], ALCDR [zhao2023ALCDR], PGPRec [yi2023PGPRec], CrossLink[huang2024CrossLink], MDGPT[yu2024MDGPT], Uni-GLM [fang2024uniglm], OMOG [liu2024OMOG], MDP-GNN [lin2025MDPGNN], UniGraph2 [he2025unigraph2], GIT [wang2025GIT], CMPGNN [wang2025bridgingCMPGNN], AEGOT-CDKT [wu2025crossAEGOTCDKT] Flatten-based GIMLET [zhao2023gimlet], GraphTranslator [zhang2024graphtranslator], GraphGPT [tang2024graphgpt], GITA [wei2024gita], InstructGraph [wang2024instructgraph], TEA-GLM [wang2024TeaGLM], GOFA [kong2025gofa], HiGPT [tang2024higpt], GraphCLIP [zhu2024graphclip], GraphWiz [chen2024graphwiz]
@@
Motivations. Although CDGL has received increasing attention, this rapidly expanding field still lacks a systematic and comprehensive review. Concretely, some existing surveys have focused on graph domain adaptation or domain generalization methods, which primarily address the distribution shift between source and target domains [shi2024GDomainAdaptationSurvey1, wu2024GDomainAdaptationSurvey2, cai2024GDomainAdaptationSurvey3]. These approaches often assume that the domains share the same feature space, meaning the node features are dimensionally and semantically aligned. Additionally, certain CDGL methods are capable of transferring across domains, but these domains typically can be further attributed to the same high-level domain, such as cross-domain graph recommendation [zang2022relatedsurvey2, chen2024relatedsurvey1]. Beyond these methods, many recent methods have proposed novel training or evaluation strategies aiming to achieve the true cross-domain generalization required by foundation models [zhao2024GCOPE, niu2024UNPrompt, liu2024OMOG, sun2025RiemannGFM, chen2025DAGPrompT]. While each of these CDGL methods contributes valuable insights toward building real GFMs, they have largely evolved in isolation. The lack of systematic analysis among them hinders researchers from gaining a holistic understanding of the CDGL landscape and identifying future research.
To fill this gap, we aim to provide a systematic survey on CDGL, helping researchers better understand the current state of research and its challenges. We review representative CDGL works to organize our first-level taxonomy, as shown in Figure 2, which includes structure-oriented, feature-oriented, and mixture-oriented approaches. We further refine this taxonomy by introducing more granularity to the initial categories.
Contributions. The contributions of this work can be summarized in the following three aspects. (1) A structured taxonomy. A broad overview of the field is presented with a structured taxonomy that categorizes existing works into three categories (Figure 2). (2) A comprehensive review. Based on the proposed taxonomy, the current research progress of cross-domain techniques for graph learning is systematically delineated. To the best of our knowledge, this is the first comprehensive survey for CDGL. (3) Some future directions. We discuss the remaining limitations of existing works and point out possible future directions.
@@ II-A Notations
Consider a graph instance represented as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{N_{v}}\}$ is the set of $N_{v}$ nodes. The set $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ defines the connections between nodes. Associated with each node $v_{i}$ is a feature vector $x_{i}\in\mathbb{R}^{D}$, and the node feature matrix is denoted as $\mathcal{X}\in\mathbb{R}^{N_{v}\times D}$. To describe the connectivity within the graph, we use the adjacency matrix $\mathcal{A}\in\{0,1\}^{N_{v}\times N_{v}}$, where the matrix entry $\mathcal{A}_{i,j}$ is defined as 1 if and only if there is an edge $(v_{i},v_{j})\in\mathcal{E}$. In this paper, the graphs from the same domain lie in the same graph domain space, denoted as $\mathcal{D}_{G}=\{(\mathcal{G}_{i},\mathcal{L}_{i})\}_{i=1}^{N_{G}}$, where each $\mathcal{G}_{i}$ is a graph, $\mathcal{L}_{i}$ denotes the corresponding label and $N_{G}$ is the number of graphs.
@@ II-B Graph Representation Learning
Shallow Embedding Methods. Shallow embedding methods, such as node2vec [grover2016node2vec] and DeepWalk [perozzi2014deepwalk], aim to map nodes into low-dimensional embeddings by preserving the network’s similarity structure. These methods utilize techniques like factorization-based methods or random walks to achieve this goal. While providing flexibility for various downstream tasks, shallow embedding methods face limitations: they cannot generate embeddings for unseen nodes and lack the capability to incorporate node features effectively.
Deep Graph Neural Networks (GNNs). Deep GNNs, including Graph Convolutional Networks (GCN) [kipf2016gcn] and Graph Attention Networks (GAT) [velivckovic2018gat], maintain input node features as constants and optimize model parameters for specific tasks. These networks typically employ a message-passing mechanism, where nodes iteratively exchange and aggregate information with their neighbors. This process leads to more expressive graph representations that can effectively handle more complex graph-based tasks. In detail, the forward process of the message-passing mechanism can be defined as:
$h_{i}^{(l)}=\mathbf{U}\left(h_{i}^{(l-1)},f_{MP}(\{h_{i}^{(l-1)},h_{j}^{(l-1)}\mid v_{j}\in\mathcal{N}_{i}\})\right)$
where $h_{i}^{(l)}$ is the feature vector of node $v_{i}$ in the $l$-th layer, and $\mathcal{N}_{i}$ is a set of neighbor nodes of node $v_{i}$. $f_{MP}$ denotes the message passing function of aggregating neighbor information, $\mathbf{U}$ denotes the update function with central node feature and neighbor node features as input. By stacking multiple layers, GNNs can aggregate messages from higher-order neighbors.
@@ II-C Cross-domain Learning
Definitions. Given $m$ source domains $\mathbb{D}_{S}=\{\mathcal{D}_{S_{1}},\dots,\mathcal{D}_{S_{m}}\}$ and $n$ target domains $\mathbb{D}_{T}=\{\mathcal{D}_{T_{1}},\dots,\mathcal{D}_{T_{n}}\}$, the objective of cross-domain learning is to learn a shared embedding space from the data in $\mathbb{D}_{S}$, and then generalize this learned embedding space to the target domains for various downstream tasks. Each domain $\mathcal{D}$ is defined as a dataset of labeled samples $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{D}}$, where $\mathbf{x}_{i}$ denotes an input sample, $y_{i}$ is the corresponding label, and $N_{D}$ is the number of samples in the domain. Each dataset is referred to as a distinct domain, where a domain is characterized by its own data distribution and task-specific context.
First, we define the model $\mathcal{M}$ as a source encoder that maps each sample $\mathbf{x}_{S_{i}}\in\mathcal{D}_{S_{i}}\subseteq\mathbb{D}_{S}$ to a unified embedding space $E$, which can be also generalized to target domains:
$\mathcal{M}:\mathbf{x}_{S_{i}}\rightarrow\mathbf{e}\in E$ (1)
[…]
The objective function for training $\mathcal{M}$ can be defined as:
$\min_{\mathcal{M}}\sum_{i=1}^{m}\mathbb{E}_{(\mathbf{x},y)\sim P_{S_{i}}(\mathbf{x},y)}\left[L_{\mathcal{M}}(\mathcal{M}(\mathbf{x}),y)\right],$ (3)
where $L_{\mathcal{M}}$ can be an unsupervised or supervised loss function (e.g., contrastive loss or classification loss), and $P_{S_{i}}(\mathbf{x},y)$ represents the joint distribution of the source domain data.
@@
$\min_{f_{T_{j}}}\sum_{j=1}^{n}\mathbb{E}_{(\mathbf{x},y)\sim P_{T_{j}}(\mathbf{x},y)}\left[L_{T}(f_{T_{j}}(\mathcal{M}(\mathbf{x})),y)\right],$ (4)
@@
[…]
@@
    NLP. The development of cross-domain learning in NLP has progressed significantly, driven by innovations in model architectures, training strategies, and data availability. Early efforts relied on domain adaptation techniques, such as feature alignment and adversarial learning, to transfer knowledge from resource-rich domains to low-resource ones. Alongside these, domain generalization techniques, which aim to improve model robustness by learning domain-invariant representations, have also been widely explored. The rise of pre-trained language models (PLMs), like BERT [devlin2019bert] and GPT [floridi2020gpt3], enabled generalization across domains by leveraging contextual embeddings and task-specific fine-tuning. More recently, large language models (LLMs), such as GPT-3 [floridi2020gpt3] and GPT-4 [achiam2023gpt4], have demonstrated strong cross-domain capabilities through massive pretraining on diverse corpora, excelling in zero-shot and few-shot learning. Furthermore, multi-modal techniques have extended cross-domain learning beyond text [zhuminigptMultimodalGPT1, chen2024mllmMultimodalGPT2], allowing models to integrate information from other modalities, such as images and structured data, to address complex tasks like visual question answering [guo2023visualquestion1] and multi-modal dialogue [wu2024visionllmMultimodalDialogue].
@@
    CV. The development of cross-domain learning in computer vision (CV) follows a similar trajectory to that of NLP [xu2025deepSurveyCV1, wang2023visionfoundation1, liu2024vmambavisionfoundation2]. Early approaches rely on domain adaptation techniques [venkateswara2017DomainAdaptationCV1] such as adversarial training, feature alignment, and style transfer to address challenges posed by domains with different visual characteristics. As the field advances, domain generalization methods [wang2022DomainGeneralizationCV1], including invariant feature learning and meta-learning, emerge to improve model robustness to unseen domains. The introduction of pre-trained vision models like ResNet [he2016deepResNet] and Vision Transformers (ViTs) [dosovitskiy2020ViT] further the progress by enabling models to generalize across domains through the reuse of learned features, which can then be fine-tuned for downstream tasks. Recently, multi-modal models like CLIP [zhu2024graphclip] and Sora [liu2024sora] demonstrate exceptional cross-domain capabilities by integrating visual and language understanding. For instance, Sora excels in tasks like image captioning, visual question answering, and zero-shot reasoning, benefiting from large-scale pretraining on multi-modal datasets. Just as in NLP, the combination of visual and textual data allows these models to achieve superior cross-domain performance [xuadvancingMultimodalCrossdomain1, xiang2025MultimodalCrossdomain2].